OpenStudy (anonymous):

Revenue is given by R(q)=550q and cost is given by C(q)=10000+5q^2 A)At what quantity is profit maximized? B)What is the total profit at this production level? Give your answer without commas.

OpenStudy (anonymous):

We wish to maximize: \[ R-C=P \]Where \(P\) is profit and all functions are defined on \(q\). So, we have: \[ \max_{q\in \mathbb{R}}\left(550q-10000-5q^2\right) \]So, we take the derivative to find the vertex: \[ \frac{d}{dq}\left(550q-10000-5q^2\right)=550-10q=0\implies\\ q=55 \]Hence we are done. For (B), simply evaluate \(P(q)\) as defined above.

OpenStudy (anonymous):

thanx man really appreciate it can u help me wit B) im on a timed test

OpenStudy (anonymous):

Just do: \[ R(55)-C(55) \]